4. Find the area of the region common to the two regions bounded by the following curves p= -6 cos(0), r = 2 – 2 cos(0)
Answer #1


Given Cardioid r= 2-2 cos 8, circle r=-6cos. 2-2 cos 8 = -6 cosa. >>4cos A = -2 27 cos=- o= 2 3 21 The given curves intersecttheta=2pi/3 y-axis 8--6cos theat r=2-2 costheta 0 x-axis -2 theta=-2pi/3